OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 2y = 12 A. (4, 0) B. (4, -6) C. (0, 6) D. (0, -6)

OpenStudy (anonymous):

why you don't try to answer questions yourself?

OpenStudy (amistre64):

the yint always starts with a 0

OpenStudy (anonymous):

Put it into slope-intercept form. Y=mx+b

OpenStudy (amistre64):

\[ax+by=c\] \[\frac{ax}{c}+\frac{by}{c}=\frac{c}{c}\] \[\frac{x}{c/a}+\frac{y}{c/b}=1\] xint = c/a yint = c/b

OpenStudy (amistre64):

in your case, 12/-2